((QUESTION 1))

The driver of a car sets the cruise control and ties the steering wheel so that the car travels at a uniform speed of 15m/s in a circle with a diameter of 100m .

Part A

Through what angular distance does the car move in 3.70min ?

Express your answer using two significant figures.

? = ____________   rad  

Part B

What arc length does it travel in this time?

Express your answer using two significant figures.

s = _________________   m  

((QUESTION 2)))

A wheel of radius 2.0m rotates at a uniform speed.

Part A

If a point on the rim of the wheel has a centripetal acceleration of 3.3m/s2 , what is the point
Answer #1

Through what angular distance does the car move in 3.70min ?
(v/r)*t=0.3*3.7*60=66.6radians
in two significant figures=67 radians

What arc length does it travel in this time?


theta*radius=66.6*50=3330m

in two significant figures=3300m

A wheel of radius 2.0m rotates at a uniform speed.

Part A

If a point on the rim of the wheel has a centripetal acceleration of 3.3m/s2 , what is the point
Answer #2

V=15 m/s

R=50m

angular velocity=15/50

=3/10 s^-1

a)angular distance=angular velocity*time

=3*3.7*60/10

=66.6 radians

b)arc lenth=angular distance*R

=50*66.6

=3330 m

c)centripetal acceleration=(tangential speed)^2 /R

tangential speed=sq rt (50*3.3)

=12.845 m/s
Answer #3

Answered Clearly , Plzz do Rate :)